In case of drawing a card, what is the probability of getting:

a. Black even number card

b. Red face

c. Black heart


1
Expert's answer
2021-01-19T15:26:18-0500

A deck of card contain "52" cards. Let "A,B,C" be three events such that ,

"A=" getting black even number card

"B=" getting red face card

"C=" getting black heart

(a) The total number of black even number cards are "10."

Therefore probability of getting black even number card is "P(A)=\frac {10}{52}=\frac{5}{26}".

(b) There are "12" face cards in a deck of "52" cards out of which there are "6" red face cards.

Therefore probability of getting a red face card is "P(B)=\frac{6}{52}=\frac{3}{26}"

(c) There are no Black heart card in a deck of 52 cards.

Therefore probability of getting black heart is "P(C)=0"